OpenStudy (anonymous):

Short answer please Find the following product: -3z(z^3 - 3z^2 + 6)

OpenStudy (anonymous):

-3z^4+9z^3-18z

OpenStudy (anonymous):

how did you get it though

OpenStudy (anonymous):

Multiply -3z with each term inside the bracket.

OpenStudy (anonymous):

okay -3z needs to times all the parts inside the brackets -3z*z^3-3z*-3z^2-3z*6 = -3z^(1+3)+9z^(1+2)-18z = -3z^4+9z^3-18z

OpenStudy (anonymous):

-3z(z^3 - 3z^2 + 6) =-3z^4+9z^3-18z

OpenStudy (anonymous):

yes correct.

OpenStudy (anonymous):

Here is a good link to use Sinz http://www.wolframalpha.com/input/?i=-3z%28z%5E3+-+3z%5E2+%2B+6%29 just type what you put here and it will figure it out for you. It also will give instructions. Also there is a key word here that help you know what you have to do and that word is "product". Product means to multiply.

OpenStudy (anonymous):

Thank you tell me if this is correct then (x+6)^2 = x^2+36

OpenStudy (anonymous):

wolfram alpha is cool

OpenStudy (anonymous):

close (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

you get the answer within seconds.

OpenStudy (anonymous):

you just missed the middle term

OpenStudy (anonymous):

nope its x^2+12x+36

OpenStudy (anonymous):

so it is x^2 + 6x + 36 right?

OpenStudy (anonymous):

okay thank you guys

OpenStudy (anonymous):

welcome